# Yongjiang Wua, Yongtao Lia, Lihua Fenga, Jiuqiang Liub, Guihai Yuc, [^2] Corresponding author

### v2

@@ Abstract
The well-known Erdős–Ko–Rado theorem states that for $n>2k$, every intersecting family of $k$-sets of $[n]:=\{1,\ldots,n\}$ has at most (${n-1\choose k-1}$ sets, and the extremal family consists of all $k$-sets containing a fixed element (called a full star). The Hilton–Milner theorem provides a stability result by determining the maximum size of a uniform intersecting family that is not a subfamily of a full star. The further stabilities were studied by Han and Kohayakawa (2017) and Huang and Peng (2024). Two families $\mathcal{F}$ and $\mathcal{G}$ are called cross-intersecting if for every $F\in\mathcal{F}$ and $G\in\mathcal{G}$, the intersection $F\cap G$ is non-empty. Let $k\geq 1,t\geq 0$ and $n\geq 2k+t$ be integers. Frankl (2016) proved that if $\mathcal{F}\subseteq\binom{[n]}{k+t}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ are cross-intersecting families, and $\mathcal{F}$ is non-empty and $(t+1)$-intersecting, then $|\mathcal{F}|+|\mathcal{G}|\leq\binom{n}{k}-\binom{n-k-t}{k}+1$. Recently, Wu (2023) sharpened Frankl’s result by establishing a stability variant. The aim of this paper is two-fold. Inspired by the above results, we first prove a further stability variant that generalizes both Frankl’s result and Wu’s result. Secondly, as an interesting application, we illustrate that the aforementioned results on cross-intersecting families could be used to establish the stability results of the Erdős–Ko–Rado theorem. More precisely, we present new short proofs of the Hilton–Milner theorem, the Han–Kohayakawa theorem and the Huang–Peng theorem. Our arguments are more straightforward, and it may be of independent interest.
@@ 1 Introduction
For any two integers $a\leq b$, let $[a,b]=\{a,a+1,\ldots,b\}$ and simply let $[n]=[1,n]$. Let $2^{[n]}$ denote the power set of $[n]$. For any $0\leq k\leq n$, let $\binom{[n]}{k}$ denote the collection of all its $k$-element subsets. A family $\mathcal{F}\subseteq 2^{[n]}$ is called $k$-uniform if $\mathcal{F}\subseteq\binom{[n]}{k}$. A family $\mathcal{F}\subseteq 2^{[n]}$ is called $t$-intersecting if $|F\cap F^{\prime}|\geq t$ for all $F,F^{\prime}\in\mathcal{F}$. If $t=1$, $\mathcal{F}$ is simply called intersecting. Two families $\mathcal{F},\mathcal{G}\subseteq 2^{[n]}$ are said to be isomorphic if there exists a permutation $\sigma$ on $[n]$ such that $\mathcal{G}=\left\{\{\sigma(x):x\in F\}:F\in\mathcal{F}\right\}$, and we write $\mathcal{F}\cong\mathcal{G}$.
@@
For $n>2k$, the equality holds if and only if =$\mathcal{F}=\big{\{}F\in\binom{[n]}{k}:x\in F\big{\}}$ for some $x\in[n]$. There are many different proofs and methods for proving the Erdős–Ko–Rado theorem; see, e.g., the Katona cycle method [30], the probabilistic method [1], the algebraic methods [20, 27, 35] and other combinatorial methods [6, 13, 28, 29, 33]. For more results on extremal set theory, we refer the interested readers to the comprehensive surveys [17, 7].
@@ 1.1 Cross-intersecting families
Let $k\geq 1$ and $n\geq 2k$ be positive integers. Let $\mathcal{F}\subseteq\binom{[n]}{k}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ be non-empty cross-intersecting families. Suppose that $|\mathcal{G}|\geq|\mathcal{F}|$. Then
@@
This result initiated the study of finding the maximum of the sum of sizes of cross-intersecting families. There are many generalizations of Hilton–Milner’s result in the past years; see, e.g., [9, 11, 15, 16, 18, 22, 38, 39]. In particular, Frankl [9] gave the following inequality, which was applied to establish the stability result of Katona theorem [29].
@@
Recently, the second author and Wu [34] sharpened Theorem 1.2 in the case $t=1$ during the study on stabilities for non-uniform $t$-intersecting families. Subsequently, Wu [40] showed the following more general extension under the constraint $|\mathcal{F}|\geq 2$.
@@
Inspired by Theorems 1.2 and 1.3, we investigate the further stability for family $\mathcal{F}$ with $|\mathcal{F}|\geq 3$. The main result of our paper is listed as below.
@@
Organization. The proof of Theorem 1.4 will be provided in Section 2. Before showing the proof, we would like to illustrate the significance of Theorem 1.4 for the stability results of the Erdős–Ko–Rado theorem. More precisely, we show that the above results on cross-intersecting families can be applied to prove the celebrated Hilton–Milner theorem, the Han–Kohayakawa theorem as well as the Huang–Peng theorem; see the forthcoming Theorems 1.5, 1.6 and 1.7, respectively. The second main contribution of this paper is to provide new short methods to prove these stability results. The detailed discussions will be presented in Section 3. To ensure our presentation clearly, some tedious arguments will be postponed to the Appendix.
1.2 Applications: Stabilities for Erdős–Ko–Rado’s theorem
In this section, we show that our result is extremely useful for the study of stabilities of $k$-uniform intersecting families. Indeed, we will provide a simple unified approach to proving several important stability results for the Erdős–Ko–Rado theorem; see, e.g., [25, 23, 26].
A family $\mathcal{F}$ that satisfies =$\mathcal{F}=\big{\{}F\in\binom{[n]}{k}:x\in F\big{\}}$ is called a full star centered at $x$, and $\mathcal{F}$ is called EKR or a star if $\mathcal{F}$ is contained in a full star. Such family is a ”trivial” example of intersecting family. If $\mathcal{F}$ is not EKR, then $\cap_{F\in\mathcal{F}}F=\emptyset$ and $\mathcal{F}$ is called non-trivial.
@@ 1.2 Applications: Stabilities for Erdős–Ko–Rado’s theorem
Let $k\geq 3$ and $n\geq 2k+1$ be integers. Let $i\in[1,k-1]$. For any $E\in\binom{[n]}{k-1}$, any $J\in\binom{[n]\backslash E}{i+1}$, $x_{0}\in J$ and $J_{i}=J\backslash\left\{x_{0}\right\}$, we define
$\mathcal{J}_{i}(n,k)=\left\{F\in\binom{[n]}{k}:x_{0}\in F,F\cap(E\cup\{j\})\neq\emptyset\text{ for each }j\in J_{i}\right\}\cup\left\{E\cup\{j\}:j\in J_{i}\right\}.$
Furthermore, let $i\in[2,k]$. For any $E\in\binom{[n]}{i}$ and $x_{0}\in[n]\backslash E$, we define
@@
To state their result, we define the family $\mathcal{K}_{2}(n,k)$ as follows. For any $E_{1},E_{2}\in\binom{[n]}{k}$ with $|E_{1}\cap E_{2}|=k-2$, and $x_{0}\in[n]\backslash(E_{1}\cup E_{2})$, let
@@
Apart from the above stabilities, there are a large number of results involving the stabilities of $t$-intersecting families; see, e.g., [2, 4, 5, 21, 34, 37] and references therein. As mentioned before, various methods involving the celebrated Erdős–Ko–Rado theorem are presented in the literature. Motivated by these methods, there are also many alternative proofs for the Hilton–Milner theorem; see, e.g., [3, 10, 12, 15, 28, 33]. For the Han–Kohayakawa theorem, Kupavskii [32] gave an alternative proof only in the case $k\geq 5$. Unfortunately, the cases $k=3$ and $k=4$ can not be proven by following his method in a straightforward way. The arguments in these cases are relatively more complicated since some exceptional extremal families are appeared. Very recently, Ge, Xu and Zhao [22] provided an alternative proof using a robust linear algebra method, and they also obtained the $t$-th level stability result for the Erdős–Ko–Rado theorem. It seems that the original proofs of Han and Kohayakawa [23] and Huang and Peng [26] are quite technical. Therefore, it is interesting to give new short proofs of the Han–Kohayakawa theorem as well as the Huang–Peng theorem.
The second goal of this paper is to show the applications of aforementioned results involving cross-intersecting families to the stability results for the Erdős–Ko–Rado theorem. We will give short proofs for the Hilton–Milner theorem, the Han–Kohayakawa theorem and the Huang–Peng theorem in a unified way. Our approach is based on applying Theorems 1.2, 1.3 and 1.4, respectively. The main purpose of this paper is to present a unified simple approach to proving the stability results for intersecting families, so we will not completely expand on these full arguments in the case $k=3$ or $4$. For interested readers, we will provide detailed proofs in the Appendix using our unified framework. In addition, it seems feasible to further extend our result on cross-intersecting families to give a short proof of the stability result in [22], we do not realize the details here, since the argument for small integer $k$ is a little bit more involved.
@@
Obviously, we have $\left|s_{i,j}(F)\right|=|F|$ and $\left|s_{i,j}(\mathcal{F})\right|=|\mathcal{F}|$. A frequently used property is that $s_{i,j}$ maintains the $t$-intersecting property of a family. A family $\mathcal{F}\subseteq 2^{[n]}$ is called shifted if for all $F\in\mathcal{F}$, $i<j$ with $i\notin F$ and $j\in F$, then $(F\backslash\{j\})\cup\{i\}\in\mathcal{F}$. It is well-known that every intersecting family can be transformed to a shifted intersecting family by applying shifting operations repeatedly. There are many nice properties for shifted families. For example, if $\mathcal{F}$ is a shifted family and $\{a_{1},\ldots,a_{k}\}\in\mathcal{F}$ with $a_{1}<\cdots<a_{k}$, then for any set $\{b_{1},\ldots,b_{k}\}$ with $b_{1}<\cdots<b_{k}$ and $b_{i}\leq a_{i}$ for each $i\in[1,k]$, we have $\{b_{1},\ldots,b_{k}\}\in\mathcal{F}$. For $\mathcal{F}\subseteq 2^{[n]}$ and $i\in[n]$, we denote $\mathcal{F}(\bar{i})=\left\{F\in\mathcal{F}:i\notin F\right\}$ and $\mathcal{F}(i)=\left\{F\backslash\{i\}:i\in F\in\mathcal{F}\right\}$.
@@
Let $k\geq 1,t\geq 0$ be integers. Let $\mathcal{F}\subseteq 2^{[n]}$ be a shifted $(t+1)$-intersecting family. Then $\mathcal{F}(\bar{1})$ is $(t+2)$-intersecting. Moreover, if $\mathcal{F}\subseteq\binom{[n]}{k+t}$ is a shifted $(t+1)$-intersecting family and $n\geq 2k+t$, then $\mathcal{F}(n)$ is $(t+1)$-intersecting.
@@
We may assume that $\mathcal{F}(\bar{1})\neq\emptyset$. For any $F_{1},F_{2}\in\mathcal{F}(\bar{1})$, let $j\in F_{1}\cap F_{2}$. By shiftedness, we have $(F_{1}\backslash\{j\})\cup\{1\}\in\mathcal{F}$. It follows from the $(t+1)$-intersecting property of $\mathcal{F}$ that $|F_{1}\cap F_{2}|=|\left((F_{1}\backslash\{j\})\cup\{1\}\right)\cap F_{2}|+1\geq t+2$. In addition, we may assume that $\mathcal{F}(n)\neq\emptyset$. For any $E_{1},E_{2}\in\mathcal{F}(n)$, we have $E_{1}\cup\{n\},E_{2}\cup\{n\}\in\mathcal{F}$. Observe that $|E_{1}\cup E_{2}\cup\{n\}|\leq 2(k+t)-(t+1)\leq n-1.$ So there exists $x\notin E_{1}\cup E_{2}\cup\{n\}$ such that $E_{1}\cup\{x\}\in\mathcal{F}$. It is immediate that $|E_{1}\cap E_{2}|=|(E_{1}\cup\{x\})\cap(E_{2}\cup\{n\})|\geq t+1,$ as desired. ∎
@@
Let $t$ be an non-negative integer. Let $k\geq 1$ and $n\geq 2k+t$ be positive integers. Let $\mathcal{F}\subseteq\binom{[n]}{k+t}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ be shifted cross-intersecting families. Then $\mathcal{F}(n)$ and $\mathcal{G}(n)$ are cross-intersecting.
@@
Next we define the lexicographic order on the $k$-element subsets of $[n]$. We say that $F$ is smaller than $G$ in the lexicographic order if $\min(F\backslash G)<\min(G\backslash F)$ holds. For $0\leq m\leq\binom{n}{k}$, let $\mathcal{L}(n,k,m)$ be the family of the first $m$ $k$-sets in the lexicographic order. The following lemma will be used in the next section, for convenience, we list it below; see [14, p.266] for a detailed proof.
@@
Proof of Theorem 1.4. For fixed $k$ and $t$, we apply induction on $n\geq 2k+t$. First let us consider the base case $n=2k+t$. For any $F\in\binom{[2k+t]}{k+t}$, the cross-intersecting property of $\mathcal{F}$ and $\mathcal{G}$ implies that $F\notin\mathcal{F}$ or $[2k+t]\backslash F\notin\mathcal{G}$. It follows that $|\mathcal{F}|+|\mathcal{G}|\leq\binom{2k+t}{k}$, as desired.
Now suppose that $n\geq 2k+t+1$ and the result holds for integers less than $n$ and fixed $k$ and $t$. By Lemma 2.3, we may assume that $\mathcal{F}$ and $\mathcal{G}$ are shifted. It follows from Lemma 2.2 that $|\mathcal{F}(\bar{n})|\geq 3$. Clearly, $\mathcal{F}(\bar{n})$ is $(t+1)$-intersecting. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(\bar{n})$ are cross-intersecting. By induction hypothesis, we get
@@
We proceed the proof by considering the following two cases.
@@
Notice that $|\mathcal{G}_{\sigma(1)}|=\binom{n-k-t-1}{k-1}$ and $|\mathcal{G}_{\sigma(2)}|\geq\binom{n-k-t-2}{k-2}$. In addition, there exist $x_{1}\in\mathcal{F}_{\sigma(1)}$ and $x_{2}\in\mathcal{F}_{\sigma(2)}$ such that $x_{1},x_{2}\notin F_{\sigma(3)}$. If $x_{1}\neq x_{2}$, then $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-3}{k-3}$; if $x_{1}=x_{2}$, then $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-2}{k-2}$. So $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-3}{k-3}$ holds in all cases. Since $\mathcal{G}_{\sigma(i)}\cap\mathcal{G}_{{\sigma(j)}}=\emptyset$ for any $i\neq j$ and $\mathcal{F}=\mathcal{F}(\bar{n})$ and $\mathcal{G}(n)$ are cross-intersecting, we have

### v3

@@ Abstract
The well-known Erdős–Ko–Rado theorem states that for $n>2k$, every intersecting family of $k$-sets of $[n]:=\{1,\ldots,n\}$ has at most (${n-1\choose k-1}$ sets, and the extremal family consists of all $k$-sets containing a fixed element (called a full star). The Hilton–Milner theorem provides a stability result by determining the maximum size of a uniform intersecting family that is not a subfamily of a full star. Further stability results were studied by Han and Kohayakawa (2017) and Huang and Peng (2024). Two families $\mathcal{F}$ and $\mathcal{G}$ are called cross-intersecting if for every $F\in\mathcal{F}$ and $G\in\mathcal{G}$, the intersection $F\cap G$ is non-empty. Let $k\geq 1,t\geq 0$ and $n\geq 2k+t$ be integers. Frankl (2016) proved that if $\mathcal{F}\subseteq\binom{[n]}{k+t}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ are cross-intersecting families, and $\mathcal{F}$ is non-empty and $(t+1)$-intersecting, then $|\mathcal{F}|+|\mathcal{G}|\leq\binom{n}{k}-\binom{n-k-t}{k}+1$. Recently, Wu (2023) sharpened Frankl’s result by establishing a stability variant. The aim of this paper is two-fold. Inspired by the above results, we first prove a further stability variant that generalizes both Frankl’s result and Wu’s result. Secondly, as an interesting application, we illustrate that the aforementioned results on cross-intersecting families could be used to establish the stability results of the Erdős–Ko–Rado theorem. More precisely, we present new short proofs of the Hilton–Milner theorem, the Han–Kohayakawa theorem and the Huang–Peng theorem. Our arguments are more straightforward, and it may be of independent interest.
@@ 1 Introduction
For any two integers $a\leq b$, let $[a,b]=\{a,a+1,\ldots,b\}$ and let $[n]=[1,n]$. Let $2^{[n]}$ denote the power set of $[n]$. For any $0\leq k\leq n$, let $\binom{[n]}{k}$ denote the collection of all its $k$-element subsets. A family $\mathcal{F}\subseteq 2^{[n]}$ is called $k$-uniform if $\mathcal{F}\subseteq\binom{[n]}{k}$. A family $\mathcal{F}\subseteq 2^{[n]}$ is called $t$-intersecting if $|F\cap F^{\prime}|\geq t$ for all $F,F^{\prime}\in\mathcal{F}$. If $t=1$, $\mathcal{F}$ is simply called intersecting. Two families $\mathcal{F},\mathcal{G}\subseteq 2^{[n]}$ are said to be isomorphic if there exists a permutation $\sigma$ on $[n]$ such that $\mathcal{G}=\left\{\{\sigma(x):x\in F\}:F\in\mathcal{F}\right\}$, and we write $\mathcal{F}\cong\mathcal{G}$.
@@
For $n>2k$, the equality holds if and only if =$\mathcal{F}=\big{\{}F\in\binom{[n]}{k}:x\in F\big{\}}$ for some $x\in[n]$. There are many different proofs and methods for proving the Erdős–Ko–Rado theorem; see, e.g., the Katona circle method [33], the probabilistic method [1], the algebraic methods [23, 30, 40] and other combinatorial methods [6, 14, 31, 32, 38]. For more results on extremal set theory, we refer the interested readers to the comprehensive surveys [18, 7].
@@ 1.1 Cross-intersecting families
Let $k\geq 1$ and $n\geq 2k$ be positive integers. Let $\mathcal{F}\subseteq\binom{[n]}{k}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ be non-empty cross-intersecting families. Under the condition that $|\mathcal{G}|\geq|\mathcal{F}|$, then
@@
This result initiated the study of finding the maximum of the sum of sizes of cross-intersecting families. There are many generalizations of the Hilton–Milner result on cross-intersecting families; see, e.g., [10, 12, 16, 17, 19, 25, 43, 44]. In particular, Frankl [10] gave the following inequality, which was applied to establish the stability result of Katona theorem [32].
@@
In the case $t=1$, Bulavka and Woodroofe [3] told us a variant by replacing the $2$-intersecting with a condition on the shadow of $\mathcal{F}$. In fact, their result can be deduced from Theorem 1.2. Recently, the second author and Wu [39] sharpened Theorem 1.2 in the case $|\mathcal{F}|\geq 2$ and $t=1$ and used this to establish stability results for non-uniform $t$-intersecting families. Subsequently, Wu [45] showed the following more general extension under the constraint $|\mathcal{F}|\geq 2$.
@@
Inspired by Theorems 1.2 and 1.3, we investigate the further stability for family $\mathcal{F}$ with $|\mathcal{F}|\geq 3$. The main result of our paper is presented below.
@@
In contrast to the Erdős–Ko–Rado theorem and the Hilton–Milner theorem (see Theorem 1.5), which gives an upper bound smaller than $k{n-2\choose k-2}$. This differs by an order of magnitude of $n$ with the bound (${n-1\choose k-1}$. While, all the upper bounds in Theorems 1.2, 1.3 and 1.4 have the same leading term (asymptotically) and only differ in the smaller order terms. It is interesting that this difference in smaller order terms will be enough to establish the stability results in Section 1.2.
Organization. The proof of Theorem 1.4 will be provided in Section 2. Before showing the proof, we would like to illustrate the significance of Theorem 1.4 for the stability results of the Erdős–Ko–Rado theorem in the next section. More precisely, we show that the above results on cross-intersecting families can be applied to prove the celebrated Hilton–Milner theorem, the Han–Kohayakawa theorem as well as the Huang–Peng theorem; see the forthcoming Theorems 1.5, 1.6 and 1.7, respectively. The second main contribution of this paper is to provide new methods to prove these stability results. The detailed discussions will be presented in Section 3. To ensure our presentation clearly, some tedious arguments will be postponed to the Appendix.
1.2 Applications: Stability for Erdős–Ko–Rado’s theorem
In this section, we show that our result is extremely useful for the study of stability of $k$-uniform intersecting families. Indeed, we will provide a simple unified approach to proving several important stability results for the Erdős–Ko–Rado theorem; see, e.g., [28, 26, 29].
A family $\mathcal{F}$ that satisfies =$\mathcal{F}=\big{\{}F\in\binom{[n]}{k}:x\in F\big{\}}$ is called a full star centered at $x$, and $\mathcal{F}$ is called EKR or a star if $\mathcal{F}$ is contained in a full star. Such family is a “trivial” example of intersecting family. If $\mathcal{F}$ is not EKR, then $\cap_{F\in\mathcal{F}}F=\emptyset$ and $\mathcal{F}$ is called non-trivial.
@@ 1.2 Applications: Stability for Erdős–Ko–Rado’s theorem
Let $k\geq 3$ and $n\geq 2k+1$ be integers. For $i\in[1,k-1]$, $E\in\binom{[n]}{k-1}$ and $x_{0}\in[n]\setminus E$, let $J_{i}$ be a subset of $[n]\setminus(E\cup\{x_{0}\})$ with $|J_{i}|=i$, we define
$\mathcal{J}_{i}(n,k)=\left\{F\in\binom{[n]}{k}:x_{0}\in F,F\cap(E\cup\{j\})\neq\emptyset\text{ for each }j\in J_{i}\right\}\cup\Bigl{\{}E\cup\{j\}:j\in J_{i}\Bigr{\}}.$
Note that $\mathcal{J}_{1}(n,k)\cong\mathcal{HM}(n,k)$. For $i\in[2,k]$, $E\in\binom{[n]}{i}$ and $x_{0}\in[n]\backslash E$, we define
@@
To construct a large intersecting family that is not EKR nor HM, note that $\mathcal{H}=\mathcal{HM}(n,k)$ has the property that $\mathcal{H}(\bar{1})=\{[2,k+1]\}$. As such, consider families $\mathcal{F}\subseteq\binom{[n]}{k}$ such that $|\mathcal{F}(\bar{1})|\geq 2$. Moreover, we want $\mathcal{F}(\bar{1})$ to be $\ell$-intersecting for $\ell$ as large as possible. When $\ell=k-1$, this gives arise to $\mathcal{J}_{i}(n,k)$. When $\ell=k-2$, this gives arise to $\mathcal{K}_{2}(n,k)$ where $\mathcal{K}_{2}(n,k)$ is defined as follows. For any $E_{1},E_{2}\in\binom{[n]}{k}$ with $|E_{1}\cap E_{2}|=k-2$, and $x_{0}\in[n]\backslash(E_{1}\cup E_{2})$, let
@@
Apart from the above stability results, there are a large number of results involving the stability for $t$-intersecting families; see, e.g., [2, 4, 5, 24, 39, 42] and references therein. As mentioned before, various methods involving the celebrated Erdős–Ko–Rado theorem are presented in the literature. Motivated by these methods, there are also many alternative proofs for the Hilton–Milner theorem; see, e.g., [3, 11, 13, 16, 31, 38]. For the Han–Kohayakawa theorem and Huang–Peng theorem, a more general result of Kupavskii [35, 37] can imply alternative proofs in the case $k\geq 5$. Unfortunately, the cases $k=3$ and $k=4$ cannot be proved following his method in a straightforward way. The arguments in these cases are relatively more complicated since some exceptional extremal families appear. Recently, Ge, Xu and Zhao [25] provided another alternative proof of the Han–Kohayakawa theorem by using a robust linear algebra method, and they also obtained the $t$-th level stability result for the Erdős–Ko–Rado theorem. It seems that the original proofs of Han and Kohayakawa [26] and Huang and Peng [29] are quite technical. Therefore, it is quite interesting to give new short proofs of the Han–Kohayakawa theorem as well as the Huang–Peng theorem.
The main contribution of this paper is to present a unified simple approach to proving the stability results for intersecting families. For simplicity, we will not completely expand on these full arguments in the case $k=3$ or $4$. For interested readers, we will provide detailed proofs in Appendices A and B using our unified framework. In addition, it seems feasible to further extend our result on cross-intersecting families to give a short proof of the stability result in [25], we do not realize the details here, since the argument for small integer $k$ is a little bit more involved.
@@
Obviously, we have $\left|s_{i,j}(F)\right|=|F|$ and $\left|s_{i,j}(\mathcal{F})\right|=|\mathcal{F}|$. A frequently used property is that $s_{i,j}$ maintains the $t$-intersecting property of a family. A family $\mathcal{F}\subseteq 2^{[n]}$ is called shifted if for all $F\in\mathcal{F}$, $i<j$ with $i\notin F$ and $j\in F$, then $(F\backslash\{j\})\cup\{i\}\in\mathcal{F}$. It is well-known that every intersecting family can be transformed to a shifted intersecting family by applying shifting operations repeatedly. There are many nice properties of shifted families. For example, if $\mathcal{F}$ is a shifted family and $\{a_{1},\ldots,a_{k}\}\in\mathcal{F}$ with $a_{1}<\cdots<a_{k}$, then for any set $\{b_{1},\ldots,b_{k}\}$ with $b_{1}<\cdots<b_{k}$ and $b_{i}\leq a_{i}$ for each $i\in[1,k]$, we have $\{b_{1},\ldots,b_{k}\}\in\mathcal{F}$. For $\mathcal{F}\subseteq 2^{[n]}$ and $i\in[n]$, we denote $\mathcal{F}(\bar{i})=\left\{F\in\mathcal{F}:i\notin F\right\}$ and $\mathcal{F}(i)=\left\{F\backslash\{i\}:i\in F\in\mathcal{F}\right\}$.
@@
Let $k\geq 1$ and $t\geq 0$ be integers. Let $\mathcal{F}\subseteq 2^{[n]}$ be a shifted $(t+1)$-intersecting family. Then $\mathcal{F}(\bar{1})$ is $(t+2)$-intersecting. Moreover, if $\mathcal{F}\subseteq\binom{[n]}{k+t}$ is a shifted $(t+1)$-intersecting family and $n\geq 2k+t$, then $\mathcal{F}(n)$ is $(t+1)$-intersecting.
@@
We may assume that $|\mathcal{F}(\bar{1})|\geq 2$ and $|\mathcal{F}(n)|\geq 2$. For any $F_{1},F_{2}\in\mathcal{F}(\bar{1})$, let $j\in F_{1}\cap F_{2}$. By shiftedness, we have $(F_{1}\backslash\{j\})\cup\{1\}\in\mathcal{F}$. It follows from the $(t+1)$-intersecting property of $\mathcal{F}$ that $|F_{1}\cap F_{2}|=|\left((F_{1}\backslash\{j\})\cup\{1\}\right)\cap F_{2}|+1\geq t+2$. For the second statement, we may assume that $\mathcal{F}(n)\neq\emptyset$. For any $E_{1},E_{2}\in\mathcal{F}(n)$, we have $E_{1}\cup\{n\},E_{2}\cup\{n\}\in\mathcal{F}$. Observe that $|E_{1}\cup E_{2}\cup\{n\}|\leq 2(k+t)-(t+1)\leq n-1.$ So there exists $x\notin E_{1}\cup E_{2}\cup\{n\}$ such that $E_{1}\cup\{x\}\in\mathcal{F}$. It is immediate that $|E_{1}\cap E_{2}|=|(E_{1}\cup\{x\})\cap(E_{2}\cup\{n\})|\geq t+1,$ as desired. ∎
@@
Let $t\geq 0$, $k\geq 1$ and $n\geq 2k+t$ be integers. Let $\mathcal{F}\subseteq\binom{[n]}{k+t}$ and $\mathcal{G}\subseteq\binom{[n]}{k}$ be shifted cross-intersecting families. Then $\mathcal{F}(n)$ and $\mathcal{G}(n)$ are cross-intersecting.
@@
We shall mention that the second statement in Lemma 2.1 and the result in Lemma 2.4 appear in [10] as Claims $1$ and $2$, respectively.
Next we define the lexicographic order on the $k$-element subsets of $[n]$. We say that $F$ is smaller than $G$ in the lexicographic order if $\min(F\backslash G)<\min(G\backslash F)$ holds. For $0\leq m\leq\binom{n}{k}$, let $\mathcal{L}(n,k,m)$ be the family of the first $m$ $k$-sets in the lexicographic order. The following lemma will be used in the next section, for convenience, we list it below; see [15, p.266] for a detailed proof.
@@
Proof of Theorem 1.4. For fixed $k$ and $t$, we apply induction on $n\geq 2k+t$. First let us consider the base case $n=2k+t$. For any $F\in\binom{[2k+t]}{k+t}$, the cross-intersecting property of $\mathcal{F}$ and $\mathcal{G}$ implies that $F\notin\mathcal{F}$ or $[2k+t]\backslash F\notin\mathcal{G}$. Hence, at most half of the sets of $\binom{[2k+t]}{k+t}\cup\binom{[2k+t]}{k}$ belong to $\mathcal{F}\cup\mathcal{G}$. Then $|\mathcal{F}|+|\mathcal{G}|\leq\binom{2k+t}{k}$. Note that $\binom{2k+t}{k}-\binom{2k+t-k-t}{k}-\binom{2k+t-k-t-1}{k-1}-\binom{2k+t-k-t-2}{k-2}+3=\binom{2k+t}{k}$. Thus the result holds.
Now suppose that $n\geq 2k+t+1$ and that the result holds for integers less than $n$ and fixed $k$ and $t$. By Lemma 2.3, we can apply all shifting operations to $\mathcal{F}$ and $\mathcal{G}$, the resulting families $\mathcal{F}^{\prime}$ and $\mathcal{G}^{\prime}$ are shifted and cross-intersecting. Moreover, $\mathcal{F}^{\prime}$ is $(t+1)$-intersecting. Thus we may assume that $\mathcal{F}$ and $\mathcal{G}$ are shifted. It follows from Lemma 2.2 that $|\mathcal{F}(\bar{n})|\geq 3$. Clearly, $\mathcal{F}(\bar{n})$ is $(t+1)$-intersecting. Note that $\mathcal{F}(\bar{n})$ and $\mathcal{G}(\bar{n})$ are cross-intersecting. By the induction hypothesis, we get
@@
We continue the proof by considering the following two cases.
@@
By the definition of $\mathcal{G}_{\sigma(1)}$, we have $|\mathcal{G}_{\sigma(1)}|=\binom{n-k-t-1}{k-1}$ as $F_{\sigma(1)}\in\binom{[n-1]}{k+t}$. Since $\mathcal{F}_{\sigma(1)}\neq\mathcal{F}_{\sigma(2)}$, there exists $x\in\mathcal{F}_{\sigma(1)}$ such that $x\notin\mathcal{F}_{\sigma(2)}$. Then $\left\{G\in\binom{[n-1]}{k-1}:x\in G,G\cap F_{\sigma(2)}=\emptyset\right\}\subseteq\mathcal{G}_{\sigma(2)}$ and hence $|\mathcal{G}_{\sigma(2)}|\geq\binom{n-k-t-2}{k-2}$. In addition, there exist $x_{1}\in\mathcal{F}_{\sigma(1)}$ and $x_{2}\in\mathcal{F}_{\sigma(2)}$ such that $x_{1},x_{2}\notin F_{\sigma(3)}$. If $x_{1}\neq x_{2}$, then $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-3}{k-3}$; if $x_{1}=x_{2}$, then $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-2}{k-2}$. So $|\mathcal{G}_{\sigma(3)}|\geq\binom{n-k-t-3}{k-3}$ holds in all cases. Since $\mathcal{G}_{\sigma(i)}\cap\mathcal{G}_{{\sigma(j)}}=\emptyset$ for any $i\neq j$ and $\mathcal{F}=\mathcal{F}(\bar{n})$ and $\mathcal{G}(n)$ are cross-intersecting, we have